OpenStudy (anonymous):

Evaluate 3x + 5 when x = 2.

OpenStudy (anonymous):

Plug in '2' for 'x': 3x + 5 3(2) + 5 Multiply 3 * 2 @AustinC

OpenStudy (anonymous):

6?

OpenStudy (yolo_king):

its quiet simple really

OpenStudy (anonymous):

Yes, now add 6 + 5

OpenStudy (yolo_king):

mhmmmm...

OpenStudy (anonymous):

Ikr @YOLO_KING

OpenStudy (anonymous):

Evaluate the expression 10x - 3 for the values given. 1,100,5,3

OpenStudy (anonymous):

so wait the first one is 11 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

1 question per post, close this and open a new one.

OpenStudy (anonymous):

The answer is 11 for the first one.

OpenStudy (anonymous):

3x+5 x=2 replace x with 2 3(2)+5 multiply 3*2 =6 now add 5

OpenStudy (anonymous):

now for that question, just do the same for one 10x-3 x=1 10(1)-3 multiply 10*1 =10 then subtract 3

OpenStudy (anonymous):

for 100 10x-3 x=100 10(100)-3 multiply 10*100 =1,000 now subtract 3 from 1,000